OpenStudy (lolipop_drum312):

Julio's father is 4 times as old as Julio. The sum of their ages is no less than 55. Enter an inequality that can be used to represent this situation in the first box, where x represents Julio's age. Enter the youngest age Julio can be in the second box.

jhonyy9 (jhonyy9):

father = 4*Julio Father +Julio >= 55 no less 55 so mean that equal or greater than 55

jhonyy9 (jhonyy9):

Julio"s age is note x so substitute the x there please

OpenStudy (lolipop_drum312):

mmm how so we substitute the x or do you mean input?

jhonyy9 (jhonyy9):

yes in place where i ve wrote Julio

OpenStudy (lolipop_drum312):

oh ok so we just write >

jhonyy9 (jhonyy9):

not just > bc this mean just greater but there in the text of your exercise ve wrote that is no less than 55 so this mean that can being equal 55 too

OpenStudy (lolipop_drum312):

oh ok so we find whats lesser or equal to 55 and input it in?

jhonyy9 (jhonyy9):

not lesser this sign mean greater or equal 55

jhonyy9 (jhonyy9):

just bc. than you read the text of your exercise there ve wrote not less than 55 ok. ?

OpenStudy (lolipop_drum312):

oh ok thanx

OpenStudy (lolipop_drum312):

wait would that be 56 or 77 idk i'm terrible at this

jhonyy9 (jhonyy9):

np yw and anytime hope helped

jhonyy9 (jhonyy9):

father age is = 4 Julio age father age = 4x so you know that 4x +x >=55 5x >= 55 divide both sides by 5 so what will get ?

OpenStudy (lolipop_drum312):

5 divided by 5 we get 1 ami doing this right

jhonyy9 (jhonyy9):

yes on the left part but on the right part what will get ?

jhonyy9 (jhonyy9):

55/5 = ?

OpenStudy (lolipop_drum312):

11

jhonyy9 (jhonyy9):

yes sure

OpenStudy (lolipop_drum312):

so 11 is the answer

jhonyy9 (jhonyy9):

so now you know that x>= 11 than how many is the age of Julio and how many is the age of father ?

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (lolipop_drum312):

yes

jhonyy9 (jhonyy9):

so than you suppose that Julio has 11 old years so this mean that father"s age is 4*11 = 44 years old ok. ?

OpenStudy (lolipop_drum312):

wai nevermind 11's the answer then 44 we were just trying to figure out the father's age right? sorry.

jhonyy9 (jhonyy9):

in the first box you need entering this inequality so in the second box the youngest Julio age

jhonyy9 (jhonyy9):

yes sure this above wrote you are right sure so now i see it that you ve understood it right sure yes ?

OpenStudy (lolipop_drum312):

yes

jhonyy9 (jhonyy9):

ok. good luck was my pleasure and anytime bye bye

OpenStudy (lolipop_drum312):

BYE